i) Given the following demand and supply functions determine the equilibrium price and quantity

Qd = 14-2P; Qs= 2+4P

ii) What factors cause the occurrence of an inverse demand curve
1
Expert's answer
2020-11-24T06:21:05-0500

(i) in equilibrium,

"Q_{d} = Q_{s}"

"14-2P=2+4P"

"6P=12"

"P=2"

"Q_{d}=Q_{s} = 14-2P=14-2(2) = 14-4 = 10"

(ii) changes in the quantity demanded lead to changes in price levels, which is the inverse of a demand curve.